1/2x - 1/4 x^2 + 4= 0
Answer by stanbon(75887)

1/2x - 1/4 x^2 + 4= 0
---------------------
Multiply thru by 4 to get:
-x^2 + 2x + 16 = 0
--------
Use the Quadratic Formula:
x = [-2 +- sqrt[4-4*-1*16]/(-2)
------
x = [-2 +- sqrt(16^2)]/(-2)
------
x = [-2 +- 16]/(-2)
----
x = -14/-2 = 7
x = (-2-16)/(-2 = 9
=====================
Cheers,
Stan H.
-----------------